OpenStudy (ineedhelpnow):

solve the system by the substitution method. X+Y=4 Y=3x

OpenStudy (amistre64):

how do you think we should go about this?

OpenStudy (amistre64):

substitution is usually quite labor intensive; but they gave you an easy one to give you the feel for it and build your confidence

OpenStudy (amistre64):

If they tell you outright that Y = some value; then wherever you encounter a Y; "substitute" that value in for it and solve

OpenStudy (ineedhelpnow):

so, x+3x=4?

OpenStudy (amistre64):

thats correct; now we solve for X to get a good solid value that we can use

OpenStudy (amistre64):

x+3x = 4 x(1+3) = 4 x(4) = 4 x = 4/4 x=1 since x=1 we know what y has to equal as well because they tell us that y = 3x y = 3(1) y= 3

OpenStudy (ineedhelpnow):

thank you!

OpenStudy (ineedhelpnow):

so, i have a similar question i need help with-- if x+3y=8 and y=2x-9 then the equation would be x+3(2x-9)=8?

OpenStudy (amistre64):

let me see x+3y = 8 y = 2x-9 ; sub this value in the top x+3(2x-9) = 8 ; and work it thru to solve for x, thats right then use your solid value for x to determine a solid value for y

OpenStudy (ineedhelpnow):

x+6x-9=8 x(1+6)-9=8 7x-9=8 7x=8+9 7x=17 x=17/7?

OpenStudy (amistre64):

you missed it up top; x +6x -27 = 8 :) the rest looks like good form tho

OpenStudy (amistre64):

in other words you 9*2 off 17 18 --- 35 so 35/7 = 5

OpenStudy (amistre64):

to work it out x+3(2x-9) = 8 x+3(2x)-3(9) = 8 x +6x -27 = 8 x(1+6) -27 = 8 x(7) -27 = 8 x(7) = 8 + 27 x = 35/7 x = 5 :)

OpenStudy (ineedhelpnow):

right! was trying to do mental math. got it! i'm also confused with x+3y=5 4x+5y=13 do i combine like terms?

OpenStudy (ineedhelpnow):

x=5-3y 4(5-3y)+5y=13 20-12y+5y=13 20-7y=13 -7y=-7 y=1 x=5-3(1) x=2

OpenStudy (amistre64):

subbing is time consuming and take alot of unneeded mathing; the next step they teach you will be "elimination" whcih takes advantage of the fact the equations can change shape and still keep the same value; kinda like fractions. Just remember that what you do to one side has to be done to it all .... x+3y = 5 ; spose we multiply this one by -4 what do we get? 4x+5y = 13 x +3y = 5 -4 ------------ -4x-12y =-20 ; this new equation has the same inherent value as the one it came from, weve only change the shape of it. Now we have something useful to us that we can use to "eliminate" one of the variables with ... watch. -4x-12y=-20 <-- our reshaped equation 4x+5y = 13 ------------ ; and add them together 0x -7y =-7 <--- now this is easily worked out for y -7y = -7 y = -7/-7 y = 1

OpenStudy (amistre64):

and since y = 1; use it in any of the equation to get the value for x x +3y = 5 x +3(1) = 5 x + 3 = 5 x = 5-3 x = 2